Calculate the molarity of 34.2 grams of sugar, C12H22O11 in 750 mL of solution.
1
Expert's answer
2020-04-22T09:20:06-0400

Molarity can be calculated using the following equation:

M = n / V = m / (Mr × V)

where M - molarity (molar concentration), n - number of moles, V - volume, m - mass, Mr - molecular weight.

Mr(C12H22O11) = (12 × 12) + (22 ×1) + (11 × 16) = 342 g/mol

750 ml = 0.75 L

As a result:

M = 34.2 g / (342g/mol × 0.75 L) = 0.13 mol/L = 0.13 M


Answer: 0.13M